# MIGHTY: Hermite Spline-based Efficient Trajectory Planning

### v2

@@ I Introduction
Trajectory planning for autonomous navigation has been extensively studied, with a wide variety of parameterizations and formulations [mellinger2011minimum, 7138978, liu2017planning, 9147300, kondo2025dynus, zhou2019robust, zhou2021ego-planner, zhou2022swarm, zhou2021raptor, tordesillas2022faster, richter2016polynomial, levy2024stitcher, wang2022geometrically, ren2025super]. For real-world deployment, fast trajectory generation is critical to ensure agile and safe flight in cluttered and dynamic environments. Approaches that use hard constraints [mellinger2011minimum, 7138978, liu2017planning, 9147300, kondo2025dynus] explicitly enforce safety and physical feasibility but often rely on computationally intensive solvers, making them difficult to deploy for high-frequency replanning on onboard processors. Soft-constraint planners such as EGO-Planner [zhou2021ego-planner], RAPTOR [zhou2021raptor], and SUPER [ren2025super] achieve faster convergence and lower computational cost. Some jointly optimize path geometry and timing within a single nonlinear formulation [wang2022geometrically, ren2025super], while others separate path generation and time allocation into two stages [richter2016polynomial, zhou2021raptor]. In general, increasing the number of decision variables, such as control points, segment durations, or derivative values, can improve performance but also enlarges the optimization problem and increases computation time. To balance efficiency and performance, recent work has explored alternative trajectory representations and optimization strategies, and building on these ideas, the following introduces our Hermite spline–based planner, MIGHTY.
@@
Polynomial [richter2016polynomial] (2016) 9th-order polynomial endpoint derivs, time decoupled limited polynomial+time space yes
Robust [zhou2019robust] (2019) non-uniform cubic B-spline B-spline cntrl pts, time decoupled limited polynomial+time space yes
EGO-Planner [zhou2021ego-planner] (2021) uniform cubic B-spline B-spline cntrl pts fixed limited polynomial+time space yes
RAPTOR [zhou2021raptor] (2021) uniform cubic B-spline B-spline cntrl pts fixed limited polynomial+time space yes
MINCO class [wang2022geometrically, zhou2022swarm, ren2025super] (2022, 2022, 2025) MINCO pos., time joint limited polynomial+time space (MINCO manifold) no
MIGHTY (proposed) Hermite Hermite cntrl pts, time. joint full polynomial+time space yes
@@ I-A Review of Unconstrained Frameworks
Differential flatness [FLIESS01061995] enables efficient optimization of quadrotor trajectories by parameterizing the flat output ($x$, $y$, $z$, yaw) and their derivatives. Given fixed timestamps at trajectory waypoints, minimum control effort (e.g., jerk/snap) problems using piecewise polynomial representation can be formulated as quadratic programs (QPs), which are convex and can be solved efficiently [mellinger2011minimum]. However, when segment durations are included as optimization variables (time allocation), the resulting joint problem becomes substantially more complex, yielding a nonlinear program (NLP) that is often ill-conditioned because of the coupling between spatial and temporal variables.
One common strategy is to decouple the problem into two stages and relax hard constraints to adopt unconstrained formulations. For example, Richter et al. [richter2016polynomial] apply piecewise 9th-order polynomials on the flat outputs with endpoint derivatives as decision variables. The trajectory optimization follows a two-stage process: an unconstrained QP minimizes snap to ensure smoothness, followed by a gradient-based time allocation step that adjusts the total trajectory duration. Collision avoidance is addressed heuristically by inserting mid-segment waypoints whenever a segment intersects obstacles, which triggers repeated QP solves until a collision-free solution is found. This approach is extended in [7353622] by improving numerical stability and incorporating additional soft state constraint costs to further ensure trajectory feasibility. Similarly, the work in [7759784] adapts this framework to achieve local real-time replanning, using endpoint derivatives as optimization variables, which minimizes snap and collisions using a potential cost function based on Euclidean signed distance field (ESDF).
@@
Another standard unconstrained framework uses B-splines for trajectory representation, with time-related variables encoded in a knot vector. B-splines offer inherent smoothness, local control, and the convex hull property, which enables efficient collision checking and enforces dynamic feasibility [9147300, zhou2019robust, zhou2021ego-planner, zhou2021raptor]. Zhou et al. [zhou2019robust] parameterize trajectories as non-uniform cubic B-splines initialized from kinodynamic A* paths. The nonlinear optimization minimizes the cost considering smoothness, a repulsive collision potential from an Euclidean distance field (EDF), and soft penalties on derivative control points exceeding maximum velocity and acceleration. Finally, an iterative time-adjustment method rescales knot spans of the non-uniform spline to guarantee dynamic feasibility. EGO-Planner [zhou2021ego-planner] and RAPTOR [zhou2021raptor] simplify the representation by using uniform B-splines with fixed knot spans, leveraging convex hull properties to enforce velocity, acceleration, and jerk constraints directly on control points. EGO-Planner optimizes trajectories via a quasi-Newton solver (L-BFGS) [liu1989lbfgs] for fast gradient-based convergence, with an additional time-reallocation and anisotropic curve-fitting refinement stage that resizes knot spans and fits a new spline to ensure feasibility and smoothness. RAPTOR also employs a two-stage process: a closed-form QP warm-start for initial smoothness, followed by a nonlinear refinement phase that penalizes collision via an ESDF and ensures dynamic limits via control point constraints.
To further improve efficiency, alternative representations of polynomial variables have been explored. Wang et al. [wang2021glst] employ an explicit diffeomorphism between coefficients and boundary derivatives, with a linear-complexity scheme for gradient evaluation towards coefficients. Built on this, MINCO [wang2022geometrically] parametrizes a segmented trajectory as piecewise polynomials, where coefficients are analytically determined by intermediate waypoints and segment durations, which are jointly optimized. A linear-complexity solver then constructs the minimum-control trajectory that minimizes the integrated control effort (e.g., jerk/snap) subject to flatness boundary conditions. Spatiotemporal deformation leverages smooth diffeomorphisms and convex constraint-elimination maps to enforce arbitrary collision and state-input constraints. However, although MINCO provides an elegant and computationally efficient formulation, since it directly maps intermediate waypoints and segment durations to polynomial coefficients using the optimality conditions of the optimal control-effort problem [wang2022geometrically], the search is, by construction, restricted to a low-dimensional manifold, potentially limiting flexibility when additional constraints or objectives are introduced. Additionally, because MINCO determines all polynomial coefficients by solving a single banded linear system subject to high-order continuity and optimality conditions, any change to one intermediate waypoint or segment duration generally changes the entire solution. In other words, MINCO does not have strict local control, meaning a small change to one waypoint or duration can propagate to many segments, complicating per-segment constraint handling for local changes. EGO-Swarm2 [zhou2022swarm] and SUPER [ren2025super] both adopt MINCO to compactly parameterize trajectories, with SUPER further optimizing exploratory and backup trajectories in each replanning cycle to balance aggressiveness and guaranteed safety. However, the intrinsic coupling within the MINCO formulation remains a limitation. MIGHTY aims to address this by optimizing over a higher-dimensional representation that preserves local control and improves flexibility for handling complex constraints.
@@ I-B MIGHTY Contributions
Summarizing the discussion above, direct unconstrained higher-order polynomial representations [richter2016polynomial] are typically too slow for real-time planning. Uniform or Non-uniform splines [zhou2019robust, zhou2021ego-planner, zhou2021raptor] can generate trajectories faster compared to spatiotemporal joint optimization; however, they can generate sub-optimal trajectories due to decoupled time adjustment. Waypoint-and-duration-based parameterizations [wang2022geometrically, zhou2022swarm, ren2025super] are restricted to a low-dimensional manifold, and a small change in one waypoint propagates globally through all polynomial coefficients. This is undesirable for real-time planning because it potentially complicates constraint/cost handling and slows down convergence.
To address the limitations of existing soft-constraint planners, MIGHTY uses a Hermite-spline representation and an unconstrained nonlinear optimizer that searches directly over spatial waypoints, endpoint derivatives, and segment durations. A Hermite spline representation also guarantees continuity without global coefficient coupling. As such, the key contributions of the paper are:
@@
    Hermite Spline Planner: Joint spatiotemporal optimization in a single solve, fully leveraging the continuous search space of Hermite splines and supporting local control of trajectory shape. The approach, MIGHTY, achieves fast solve times and superior performance.
@@
    Extensive Hardware Experiments: Long-duration flights, high-speed flights, and dynamic obstacle avoidance scenarios using a LiDAR-based perception and localization system. MIGHTY achieves collision-free long-duration flights and a top speed of $6.7\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$ in high-speed flight experiments.
@@ III-C Objective and Closed-Form Gradient
To make the formulation basis-independent, we express the objective and its gradients with respect to a generic control-point (or coefficient) vector $\mathbf{c}$, independent of the chosen parameterization (Hermite end states, Bézier control points, etc.). As a concrete example, the affine map between Hermite and Bézier on segment $s$ is
@@
And $\partial\mathbf{c}_{s}/\partial T_{s}$ used in the coefficient path in Eq. (4) can be computed in closed form as
$\frac{\mathbf{c}_{s,0}}{T_{s}}$ $=0,\quad\frac{\mathbf{c}_{s,1}}{T_{s}}=\tfrac{1}{5}\,\mathbf{v}_{s},\quad\frac{\mathbf{c}_{s,2}}{T_{s}}=\tfrac{2}{5}\,\mathbf{v}_{s}+\tfrac{T_{s}}{10}\,\mathbf{a}_{s},$ (5)
$\frac{\mathbf{c}_{s,3}}{T_{s}}$ $=-\tfrac{2}{5}\,\mathbf{v}_{s+1}+\tfrac{T_{s}}{10}\,\mathbf{a}_{s+1},\quad\frac{\mathbf{c}_{s,4}}{T_{s}}=-\tfrac{1}{5}\,\mathbf{v}_{s+1},\quad\frac{\mathbf{c}_{s,5}}{T_{s}}=0.$
@@ III-D Reparameterizations for Derivative Variables
and set $\hat{\mathbf{v}}_{i}=\bar{T}_{i}\,\mathbf{v}_{i},\quad\hat{\mathbf{a}}_{i}=\bar{T}_{i}^{2}\,\mathbf{a}_{i}.$ This reparameterization removes the explicit $1/T_{s}$ factors from the gradients and balances sensitivity across short and long segments. Our ablation study in Section IV-C confirms the expected improvement in performance.
@@ IV-A Representation Benchmarking: Simple Case
We first benchmark MIGHTY against GCOPTER (which uses MINCO [wang2022geometrically]) on a simple corner-avoidance task with fixed start and goal (Fig. 1). For a fair comparison, we implemented GCOPTER’s optimization cost in MIGHTY and use the same optimizer (L-BFGS [liu1989lbfgs]) and stopping tolerances. Both methods start from the same initial guess.
@@
The result summary in Fig. 2 shows that MIGHTY achieves a shorter travel time, with slightly faster computation and shorter path than GCOPTER. While GCOPTER produces smoother trajectories with lower jerk due to its constraint within the MINCO manifold, MIGHTY explores a larger search space to identify faster and more effective trajectories. This flexibility enables MIGHTY to achieve superior overall performance.
IV-B Representation Benchmarking: Complex Case

### v3

@@ I Introduction
Trajectory planning for autonomous navigation has been extensively studied with various parameterizations [13, 3, 10, 18, 7, 23, 25, 26, 24, 19, 17, 8, 22, 16]. Hard-constraint approaches [13, 3, 10, 18, 7] explicitly enforce safety but require computationally intensive solvers unsuitable for high-frequency replanning. Soft-constraint planners (EGO-Planner [25], RAPTOR [24], SUPER [16]) achieve faster convergence. Some jointly optimize geometry and timing [22, 16], while others decouple path and time allocation [17, 24]. Increasing decision variables improves performance but enlarges the problem. Building on these ideas, we introduce MIGHTY, a Hermite spline-based planner.
@@
Geom. Deriv.
Polynomial [17] (2016) 9th-order polynomial endpoint derivs, time decoupled limited polynomial+time space no (global) direct but global
Robust [23] (2019) non-uniform B-spline B-spline cntrl pts, time decoupled limited polynomial+time space yes indirect
EGO-Planner [25] (2021) uniform B-spline B-spline cntrl pts fixed limited polynomial+time space yes indirect
RAPTOR [24] (2021) uniform B-spline B-spline cntrl pts fixed limited polynomial+time space yes indirect
MINCO class [22, 26, 16] (2022, 2022, 2025) MINCO pos., time joint limited polynomial+time space (MINCO class) no (global) indirect
MIGHTY (proposed) Hermite spline Hermite cntrl pts, time. joint full polynomial+time space yes direct & local
@@ I-A Review of Unconstrained Frameworks
Differential flatness [5] enables efficient optimization of quadrotor trajectories by parameterizing the flat output and their derivatives. Given fixed timestamps at trajectory waypoints, minimum control effort (e.g., jerk/snap) problems using piecewise polynomial representation can be formulated as quadratic programs (QPs), which are convex and can be solved efficiently [13]. However, when segment durations are included as optimization variables (time allocation), the resulting joint problem becomes substantially more complex, yielding a nonlinear program (NLP) that is often ill-conditioned because of the coupling between spatial and temporal variables.
Decoupling into two stages relaxes hard constraints for unconstrained formulations. Richter et al. [17] use 9th-order polynomials with endpoint derivatives: a QP minimizes snap, then gradient-based time allocation adjusts duration. Endpoint derivatives are decision variables, so there is direct derivative parameterization. However, the free derivatives are globally coupled, so there is no local control. Collision avoidance inserts mid-segment waypoints, triggering repeated solves. Extensions [1, 14] improve stability and enable real-time replanning via ESDF-based potential costs.
@@
Another standard unconstrained framework uses B-splines for trajectory representation, with time-related variables encoded in a knot vector. B-splines offer inherent smoothness and geometric local control: changing one control point affects only a bounded neighborhood of the curve. The convex hull property enables efficient collision checking and enforces dynamic feasibility [18, 23, 25, 24]. However, B-splines do not provide direct local control of higher-order derivatives. Derivatives at any curve point are implicit functions of multiple neighboring control points. One can bound derivatives efficiently via the convex hull property, but cannot independently prescribe a specific velocity or acceleration at a given knot through a single variable. This geometric local control property has been widely exploited for efficient trajectory refinement and replanning in cluttered environments [6, 15, 20, 4]. Zhou et al. [23] parameterize trajectories as non-uniform cubic B-splines initialized from kinodynamic A* paths. The nonlinear optimization minimizes the cost considering smoothness, a repulsive collision potential from an Euclidean distance field (EDF), and soft penalties on derivative control points exceeding maximum velocity and acceleration. Finally, an iterative time-adjustment method rescales knot spans of the non-uniform spline to guarantee dynamic feasibility. EGO-Planner [25] and RAPTOR [24] simplify the representation by using uniform B-splines with fixed knot spans, leveraging convex hull properties to enforce velocity, acceleration, and jerk constraints directly on control points. EGO-Planner optimizes trajectories via a quasi-Newton solver (L-BFGS) [9] for fast gradient-based convergence, with an additional time-reallocation and anisotropic curve-fitting refinement stage that resizes knot spans and fits a new spline to ensure feasibility and smoothness. RAPTOR also employs a two-stage process: a closed-form QP warm-start for initial smoothness, followed by a nonlinear refinement phase that penalizes collision via an ESDF and ensures dynamic limits via control point constraints.
Wang et al. [21] employ diffeomorphism between coefficients and derivatives with linear-complexity gradients. MINCO [22] builds on this: coefficients are analytically determined by waypoints and durations, jointly optimized via linear-complexity solver for minimum jerk/snap. MINCO excels when minimum control effort dominates, but searches only the MINCO class. MINCO’s waypoint positions are decision variables, but polynomial coefficients are determined through a banded linear system that couples adjacent segments, so there is no geometric local control. Higher-order derivatives are not decision variables, and they are analytically determined from waypoints and durations, so there is no direct higher-derivative control. Under added objectives, this restriction can yield suboptimal performance compared to more general representations. EGO-Swarm2 [26] and SUPER [16] adopt MINCO; SUPER adds exploratory/backup trajectories for safety.
MIGHTY optimizes positions and derivatives at each knot and time durations as explicit variables, enabling local control of waypoints and higher-order dynamics.
@@ I-B MIGHTY Contributions
Table I summarizes the discussion above. Unconstrained higher-order polynomial representations [17] are typically too slow for real-time planning. Uniform or Non-uniform splines [23, 25, 24] can generate trajectories faster compared to spatiotemporal joint optimization; however, they can generate sub-optimal trajectories due to decoupled time adjustment. Waypoint-and-duration-based parameterizations [22, 26, 16] are restricted to the low-dimensional subspace induced by the MINCO parameterization, lack direct parameterization of higher-order derivatives at individual knots, and exhibit waypoint propagation where changes to one waypoint affect multiple segments through the globally coupled coefficient system. While these effects enable compact representation and analytical optimality for minimum-control objectives, they can complicate constraint/cost handling beyond minimum-control and limit local control of trajectory shape and higher-order dynamics.
To address the limitations of existing soft-constraint planners, MIGHTY uses a Hermite-spline representation and an unconstrained nonlinear optimizer that searches directly over spatial waypoints, endpoint derivatives, and segment durations. A Hermite spline representation guarantees continuity without global coefficient coupling, which fundamentally enables direct local control over both positions and higher-order dynamics at each knot. These properties could be exploited for efficient trajectory refinement (as in B-spline methods [6, 15, 20, 4]) but are not available in MINCO’s globally coupled parameterization. The key contributions of the paper are:
@@
    Hermite Spline Planner: Joint spatiotemporal optimization in a single solve, with explicit parameterization of positions and derivatives at each knot. This representation enables local control of higher-order dynamics and trajectory shape, which the optimizer can leverage when needed. The approach, MIGHTY, achieves fast solve times and superior performance.
@@
    Extensive Hardware Experiments: Long-duration and high-speed flights, and dynamic obstacle avoidance scenarios using a LiDAR-based perception and localization system. MIGHTY achieves collision-free flights and a top speed of $6.7\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$ in high-speed flight experiments.
@@ III-C Objective and Closed-Form Gradient
To make the formulation basis-independent, we express the objective and its gradients w.r.t a generic control-point (or coefficient) vector $\mathbf{c}$, independent of the chosen parameterization (Hermite end states, Bézier control points, etc.). As a concrete example, the affine map between Hermite and Bézier on segment $s$ is
@@
And $\partial\mathbf{c}_{s}/\partial T_{s}$ ($\nabla_{T_{s}}\mathbf{c}_{s}$) used in the coefficient path in Eq. (4) can be computed in closed form as
$\nabla_{T_{s}}\mathbf{c}_{s,0}$ $=\mathbf{0},\ \nabla_{T_{s}}\mathbf{c}_{s,1}=\tfrac{1}{5}\,\mathbf{v}_{s},\ \nabla_{T_{s}}\mathbf{c}_{s,2}=\tfrac{2}{5}\,\mathbf{v}_{s}+\tfrac{T_{s}}{10}\,\mathbf{a}_{s},$ (5)
$\nabla_{T_{s}}\mathbf{c}_{s,3}$ $=-\tfrac{2}{5}\,\mathbf{v}_{s+1}+\tfrac{T_{s}}{10}\,\mathbf{a}_{s+1},\ \nabla_{T_{s}}\mathbf{c}_{s,4}=-\tfrac{1}{5}\,\mathbf{v}_{s+1},$
$\nabla_{T_{s}}\mathbf{c}_{s,5}$ $=\mathbf{0}.$
@@ III-D Reparameterizations for Derivative Variables
and set $\hat{\mathbf{v}}_{i}=\bar{T}_{i}\,\mathbf{v}_{i},\quad\hat{\mathbf{a}}_{i}=\bar{T}_{i}^{2}\,\mathbf{a}_{i}.$ This reparameterization aims to reduce the explicit $1/T_{s}$ and $1/T_{s}^{2}$ factors’ effect in the gradients w.r.t. the derivative decision variables by optimizing the scaled variables $\{\hat{\mathbf{v}}_{i},\hat{\mathbf{a}}_{i}\}$ instead of $\{\mathbf{v}_{i},\mathbf{a}_{i}\}$. At the endpoints, the scaling uses the adjacent segment duration (no averaging), while at interior knots we use the local average duration to balance the influence of the two neighboring segments; in that case, the inverse-duration effects are mitigated but cannot be canceled for both segments simultaneously unless their durations match. Our ablation study in Section IV-E confirms the expected improvement in performance.
@@ IV-A Representation Benchmarking: Simple Case
We first benchmark MIGHTY against GCOPTER (which uses MINCO [22]) on a simple corner-avoidance task with fixed start and goal (Fig. 1). For a fair comparison, we implemented GCOPTER’s optimization cost in MIGHTY and use the same optimizer (L-BFGS [9]) and stopping tolerances. Both methods start from the same initial guess.
@@
The result in Fig. 2 shows that MIGHTY achieves a shorter travel time, with slightly faster computation and shorter path than GCOPTER. While GCOPTER produces smoother trajectories with lower jerk due to its constraint within the MINCO class, MIGHTY explores a larger search space to identify faster and more effective trajectories. This flexibility enables MIGHTY to achieve superior performance, and Section IV-B further investigates the jerk-performance trade-off.
IV-B Smooth Weight Sweep Benchmarking
Fig. 2 shows MIGHTY yields higher jerk than GCOPTER while achieving better travel time. To quantify this trade-off, we sweep $w_{\mathrm{smooth}}\in\{${10}^{-5}$,\ldots,${10}^{3}$\$\} in the same scenario. Fig. 3 shows that as $w_{\mathrm{smooth}}$ increases, MIGHTY progressively reduces jerk, eventually surpassing GCOPTER’s levels of smoothness. At $w_{\mathrm{smooth}}=${10}^{2}$\$, MIGHTY matches GCOPTER’s jerk while maintaining better performance and lower computation time, demonstrating that MIGHTY’s higher jerk is a tunable design choice rather than an inherent limitation.
Figure 3: Effect of jerk-smoothness weighting in the simple corner-avoidance scenario (Fig. 1). As $w_{\mathrm{smooth}}$ increases, MIGHTY’s jerk decreases toward GCOPTER’s jerk (orange dotted reference). At $w_{\mathrm{smooth}}=${10}^{2}$\$ (vertical red line), MIGHTY matches GCOPTER’s jerk with comparable performance and lower computation time.
IV-C Local Control of Higher-Order Dynamics
MINCO-based planners lack direct parameterization of higher-order derivatives at individual knots; these quantities are derived from a globally coupled coefficient system (Sec. I-A). While GCOPTER can add derivative constraints as soft penalties, MIGHTY’s Hermite parameterization treats knot positions and velocities as explicit optimization variables. To evaluate this difference, we extend the simple corner-avoidance scenario (Fig. 1) with soft reference tracking terms for position and velocity at the penultimate knot, which is relevant for waypoint velocity specification and hover-at-goal behaviors.
Figure 4: Local control of higher-order dynamics: position and velocity reference tracking. GCOPTER and MIGHTY trajectories are color-mapped by speed. Both planners enforce a velocity reference $\mathbf{v}_{\text{ref}}=(0,0,0)\,$\mathrm{m}\mathrm{/}\mathrm{s}$\$ and position reference at the penultimate knot (one knot before the goal). As shown, MIGHTY achieves better reference tracking accuracy.
We augment the objective function with soft position and velocity reference terms
$J_{\text{ref}}=\tfrac{1}{2}w_{\mathbf{p}_{\text{ref}}}\lVert\mathbf{p}(t_{k})-\mathbf{p}_{\text{ref}}\rVert^{2}+\tfrac{1}{2}w_{\mathbf{v}_{\text{ref}}}\lVert\mathbf{v}(t_{k})-\mathbf{v}_{\text{ref}}\rVert^{2},$
where $\mathbf{p}_{\text{ref}}$ and $\mathbf{v}_{\text{ref}}$ are desired position and velocity at knot $k$, and $w_{\mathbf{p}_{\text{ref}}}$ and $w_{\mathbf{v}_{\text{ref}}}$ are penalty weights. For this benchmark, we set $k$ to the penultimate knot (one knot before the goal), $\mathbf{v}_{\text{ref}}=(0,0,0)\,$\mathrm{m}\mathrm{/}\mathrm{s}$\$ to enforce near-hover at the goal, and $\mathbf{p}_{\text{ref}}$ to the goal position. We use $w_{\mathbf{p}_{\text{ref}}}=$1\text{\times}{10}^{3}$\$ and $w_{\mathbf{v}_{\text{ref}}}=$1\text{\times}{10}^{3}$\$, with $v_{\max}=$2.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$ and $w_{\mathrm{smooth}}=${10}^{-5}$\$ for MIGHTY. All other settings match those in Sec. IV-F. We evaluate both GCOPTER and MIGHTY across 100 runs in the same simple corner-avoidance environment. Fig. 4 shows example trajectories from both planners, and Table II summarizes the results.
TABLE II: Local control of higher-order dynamics benchmark. We added position/velocity soft penalties on both GCOPTER and MIGHTY and evaluated performance. Best results are in green.
Planner $e_{\mathbf{p}_{\text{ref}}}$ [m] $e_{\mathbf{v}_{\text{ref}}}$ [m/s] $T_{\mathrm{trav}}$ [s] $T_{\mathrm{comp}}$ [ms]
GCOPTER 0.017 1.0 19.7 7.4
MIGHTY 0.0041 0.16 14.8 10.6
[…]
IV-D Representation Benchmarking: Complex Case